OpenStudy (lionhowlter):

What are the x and y intercepts of the line 2x-3y=-12? a.) x- intercept: 6, y intercept: -4 b.) x- intercept: 4, y intercept: -6 c.) x - intercept: -4, y intercept: 6 d.) x- intercept: -6 y intercept: 4

OpenStudy (lionhowlter):

help please?

OpenStudy (taemin):

2x – 3y = 12 u will get... (2/3)x - y = 4 y = (2/3)x -4 now this equation is in the form of y = mx + b where b is the y intercept... so.. in our equation y intercept is " -4" to find x intercept... u have to find the value of "x" when "y" is 0... y = (2/3)x -4 0 = (2/3)x - 4 (2/3)x = 4 2x = 12 x = 6

OpenStudy (taemin):

a.) x- intercept: 6, y intercept: -4